OpenStudy (sophiesky):

Alright, one more question. Write the standard form of the equation of the line perpendicular to the graph of the equation 3x+2y=6 through the point (1,3).

OpenStudy (sophiesky):

@jim_thompson5910 Could you help me with one last question?

jimthompson5910 (jim_thompson5910):

We're given 3x+2y=6 To find the perpendicular equation, we simply swap the '3' and '2'. Then we make the y coefficient have an opposite sign So we'll have 2x-3y = C where C is some constant

jimthompson5910 (jim_thompson5910):

plug in the given point (1,3) to get 2x-3y = C 2*(1) - 3*(3) = C C = ???

OpenStudy (sophiesky):

Why do we swap 3 and 2 and make y negative? Is it because we're looking for the perpendicular equation?

OpenStudy (sophiesky):

Why do we swap 3 and 2 and make y negative? Is it because we're looking for the perpendicular equation?

jimthompson5910 (jim_thompson5910):

Well it's a shortcut I'm using. Did you want to do this the long way?

jimthompson5910 (jim_thompson5910):

The rule I used is if you had Ax+By = C, then the perpendicular equation would be Bx-Ay = D

OpenStudy (sophiesky):

Well no... Shortcuts are greatly appreciated.. But I just want to make sure I'm completely understanding what I'm doing, for future reference. :)

jimthompson5910 (jim_thompson5910):

Recall that if you had a slope of say 5/7, then the perpendicular slope is -7/5 you flip the fraction and you flip the sign

OpenStudy (sophiesky):

Ohhh alright

jimthompson5910 (jim_thompson5910):

so A and B swapping and one changing signs is following that same idea

OpenStudy (sophiesky):

Okay so C is -7.

jimthompson5910 (jim_thompson5910):

yes, so 2x-3y = C becomes 2x-3y = -7

OpenStudy (sophiesky):

And that's the perpendicular equation in standard form?

jimthompson5910 (jim_thompson5910):

3x+2y = 6 is the original equation 2x-3y = -7 is the perpendicular equation through (1,3)

jimthompson5910 (jim_thompson5910):

`And that's the perpendicular equation in standard form?` yes

jimthompson5910 (jim_thompson5910):

the long way is to get the equation into y = mx+b form, find the perpendicular slope, then determine b, then get everything back into standard form

OpenStudy (sophiesky):

Does the flip thing you did work every time?

jimthompson5910 (jim_thompson5910):

yes it does --------------------------- Ax+By = C is equation 1 Bx-Ay = D is equation 2 I claim that equation 1 and equation 2 are perpendicular. To prove this claim, I'll need the slopes first solve Ax+By = C for y to get y = (-A/B)x + C/B The slope here is -A/B solve Bx-Ay = D for y to get y = (B/A)x - D/A the slope is B/A Multiply the two slopes -A/B and B/A and you'll get -1 Rule: if two lines have slopes that multiply to -1, then they are perpendicular lines so that proves equation 1 and equation 2 are perpendicular lines

jimthompson5910 (jim_thompson5910):

In this case, A = 3 B = 2 C = 6 D = -7

OpenStudy (sophiesky):

Awesome! Thank you so much!!!!

jimthompson5910 (jim_thompson5910):

no problem